OpenStudy (anonymous):

1/6 - 2/7 =? show your work.

OpenStudy (mathstudent55):

You need a common denominator.

OpenStudy (anonymous):

here is a site http://www.webmath.com/_answer.php

OpenStudy (anonymous):

you just type in anything it has all different equations and everything i already typed it in for you

OpenStudy (anonymous):

it says URL not found

OpenStudy (anonymous):

scroll all the way to the bottom and the answer is 5 -- 42

OpenStudy (anonymous):

the site is called webmath.com

OpenStudy (anonymous):

here i copied it down this will show you everything you need to know The problem here is to subtract and These two fractions do not have the same denominators (lower numbers), so we must first find a common denominator of the two fractions, before subtracting them. For the denominators here, the 6 and 7, a common denominator for both is 42. With the common denominator, the becomes a and the becomes a So now our subtraction problem becomes this... The problem here is to subtract and These two fractions have the same denominators, but numerator of the first fraction, the 7, is less than the numerator of the second fraction, the 12. This means, before subtracting the fractions, we have to "borrow" a 1 from the whole number 0, making it a -1. We can add this borrowed 1 to the fraction 7/42, like this: 7/42 + 1 is 7/42 + 42/42 = 49/42 After doing all of this, the becomes a , and we can now perform the subtraction... The problem here is to subtract and Since these two fractions have the same denominators (the numbers under the fraction bar), we can subtract them by simply subtracting the numerators (the 49 - 12 = 37), while keeping the same denominator (the 42). And, we just subtract the whole number parts, or -1 - 0 = -1. Our answer here is: So far, the answer to your problem is: This answer is correct and may be left as is. But, having a negative whole number is not the usual way to write down fractions. So, why don't we just convert it into an improper fraction and use that as the answer?!?! To convert to an improper fraction: multiply the 42 (the denominator), and the -1 (the whole number). To this product, add the 37 (the numerator) giving -5, to form the new numerator, and use the 42 as the new denominator. So, as an improper fraction is The final answer is:

OpenStudy (anonymous):